OpenStudy (anonymous):

Use math induction to show that (n2 - n + 2) is divisible by 2 for all natural numbers n.

OpenStudy (anonymous):

Here's the process of induction. When n=1, 1^2-1+2=2. 2 is divided by 2. When n=2, 2^2-2+2=4. 4 is divided by 2. When n=3, 3^2-3+2=8. 8 is divided by 2. Suppose when n=k, k^2-k+2 is divided by 2. Let n=k+1. Then (k+1)^2-(k+1)+2=k^2+2k+1-k-1+2 =k^2+k+2=k(k+1)+2 Note one of k and k+1 must be even because they are consecutive natural numbers. Then k(k+1)+2 will be divisible by 2. Consequently, n^2-n+2 is divided by 2 for all natural numbers n.

OpenStudy (schrodinger):

Marly, is that 1^(2-1+2) or is it 1^2-1+2?